OpenStudy (anonymous):

Let f(x) = -20x2 + 14x + 12 and g(x) = 5x - 6. Find f/g and state its domain. –4x – 2; all real numbers except 6/5 –4x – 2; all real numbers 5x - 6; all real numbers except 6/5 5x - 6; all real numbers

OpenStudy (anonymous):

\[\frac{f}{g}(x)=\frac{f(x)}{g(x)}=\frac{-20x^2+14x+12}{5x-6}\] Long division: